OpenStudy (anonymous):

Which ratio is equal to 5/6?

OpenStudy (anonymous):

10/12

OpenStudy (anonymous):

Here are the choices. A. 15/16 B. 25/36 C. 30/36 D. 60/70

OpenStudy (anonymous):

Plenty of ratios are equal to 5/6. If you multiply the numerator by any number, you can do the same to the denominator and the resulting ratio will always be equal to 5/6.

OpenStudy (anonymous):

Simply the fractions.

OpenStudy (anonymous):

15/16 can not be simplified. So that is not the answer.

OpenStudy (anonymous):

25/36 cannot be simplified either, so that is not the answer.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Yes, it is C. He simplified.

OpenStudy (anonymous):

30/36 if all divided by 6 is equal to 5/6

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Here is another one, Write a proportion for the following problem. Eric is making a scale drawing of the ground floor of his house. If the scale is 2 in = 5 ft, what number of inches will he use for a length of 55 ft?

OpenStudy (anonymous):

Here are the choices, A. 2/5 = 55/x B. 2/55 = x/5 C. 2/5 = x/55

OpenStudy (anonymous):

The answer is B.

OpenStudy (anonymous):

Not it isn't.

OpenStudy (anonymous):

I'm sorry, C.

OpenStudy (anonymous):

in/ft = in/ft

OpenStudy (anonymous):

Yes, I misread the answers.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Inches to inches and feet to feet.

OpenStudy (anonymous):

Thank you for your help.

OpenStudy (anonymous):

You're welcome.